OpenStudy (anonymous):

ANSWER PLEASE If 18 meters of high-resistance wire with a diameter of 1.2 mm has a resistance of 10 ohms, what is the resistance of 27 m of the same type of wire with a diameter of 1.5 mm?

OpenStudy (anonymous):

can somebody please help me by explaining it and show me the answer and how to get it

OpenStudy (cj7529):

the resistivity of a resistor is given by resistivity = (RA)/L where R=resistance A=Cross sectional area of wire L length of the wire The resistivity of the given type of wire can be assumed to be constant here, So \[\frac{ R1 A1 }{ L1 } = \frac{ R2 A2 }{ L2 }\] R2 is what youre after so you can rearange to get \[\frac{ R1 A2 L2 }{ L1 A2 } \] then just plug in the numbers in the question. that cool?